# Constructing Variables Using Classifiers as an Aid to Regression: An Empirical Assessment

### 11 Mar 2024 v1

@@ 5 Conclusion
This article has proposed a method for the automatic creation of variables (in the case of regression) that complement the information contained in the initial input vector, the explanatory variables. The method works as a pre-processing step in which the continuous values of the variable to be regressed are discretized into a set of intervals. These intervals are used to define threshold values. Next, classifiers are trained to predict whether the value to be regressed is less than or equal to each of these thresholds. The different outputs of the classifiers are then concatenated into an additional vector of variables that enriches the initial vector of explanatory variables native to the regression problem. The results are encouraging, even if they mainly benefit three of the five regressors for which the method has been tested. A first improvement would be to extract a more informative vector from the classifiers, such as tree leaf identifiers. A second improvement could be to design a neural architecture that groups together all the steps of the proposed method.
The proposed method also opens up certain perspectives. The attentive reader will probably have noticed that, as defined in the proposed implementation, the vector $X^{\prime}$ is in fact an estimate of the conditional cumulative density of $y$ knowing $X$. It would therefore be possible, via an expectation calculation, to dispense with regressors altogether in the case where $S$ is sufficiently high to have a fairly good estimate of this cumulative density. This last point will probably be the subject of future work.

### 13 Mar 2024 v2

@@ 5 Conclusion
This paper has proposed a method for automatically generating variables (in the case of regression) that complement the information contained in the initial input vector, the explanatory variables. The method works as a pre-processing step in which the continuous values of the variable to be regressed are discretized into a set of intervals. These intervals are used to define threshold values. Classifiers are then trained to predict whether the value to be regressed is less than or equal to each of these thresholds. The different outputs of the classifiers are then concatenated into an additional vector of variables that enriches the initial vector of explanatory variables native to the regression problem. The results are encouraging, although they mainly benefit three of the five regressors for which the method has been tested. A first improvement would be to extract a more informative vector from the classifiers, such as tree leaf identifiers. A second improvement could be to design a neural architecture that combines all the steps of the proposed method.
The proposed method also opens up certain perspectives. The attentive reader will have noticed that the vector $X^{\prime}$, as defined in the proposed implementation, is in fact an estimate of the conditional cumulative density of $y$, knowing $X$. It would therefore be possible to dispense with regressors altogether, via an expectation calculation, in the case where $S$ is sufficiently high to have a fairly good estimate of this cumulative density. This last point is likely to be the subject of future work.